OpenStudy (anonymous):

Find the area of a regular heptagon with an apothem 18.7 feet long and a side length of 18 feet.

OpenStudy (anonymous):

A = 1/2ap = 1/2(18.7)(7 * 18) = 18.7 * 63 = 1178.1 ft²

OpenStudy (anonymous):

no wait im confused

OpenStudy (anonymous):

a. 7.875 cm2 c. 31.5 cm2 b. 15.75 cm2 d. 12.6 cm2

OpenStudy (anonymous):

It isnt even using the same measurements as the question?o-o

OpenStudy (anonymous):

I think that's another question lol

OpenStudy (anonymous):

ohh haa i was looking at the next one opps

OpenStudy (anonymous):

can you help me some more?

OpenStudy (anonymous):

@Yuba

OpenStudy (anonymous):

Wish I could honestly, but Im really not good at these things lol